### v2

# CanadaFireSat: Toward high-resolution wildfire forecasting with multiple modalities

Abstract
Canada experienced in 2023 one of the most severe wildfire seasons in recent history, causing damage across ecosystems, destroying communities, and emitting large quantities of $\text{CO}_{2}$. This extreme wildfire season is symptomatic of a climate-change-induced increase in length and severity of fire seasons affecting the boreal ecosystem. Therefore, it is critical to empower wildfire management in boreal communities with better monitoring solutions. Wildfire probability maps are an important tool for understanding the likelihood of wildfire occurrence and the potential severity of future wildfires. Fire forecasting tools based on Earth observation data exist, but they are limited both by the lack of label information and by their reliance on coarse-resolution environmental drivers and satellite products, which leads to wildfire occurrence prediction of reduced resolution, typically around $\sim 0.1$°. To tackle these two limitations, this paper presents a benchmark dataset, CanadaFireSat, and baseline methods for high-resolution wildfire forecasting at $100$ m across Canada. CanadaFireSat leverages multi-modal data from high-resolution multi-spectral satellite images (Sentinel-2), mid-resolution satellite products (MODIS), and environmental factors (ERA5). We experiment with convolutional (CNN) and transformer (ViT) architectures. We observe that using multi-modal temporal inputs outperforms single-modal temporal inputs across all metrics, achieving a peak performance of $60.3\%$ in F1 score for the 2023 wildfire season, a season never seen during model training. This demonstrates the potential of multi-modal deep learning for wildfire forecasting at high-resolution and continental scale.
keywords:
Wildfire Forecasting , Benchmark Dataset , Multi-modal Learning , Deep Learning , Boreal Ecosystem
[…]
[inst1]organization=EPFL,addressline=Route des Ronquos 86, city=Sion, postcode=1950, state=Wallis, country=Switzerland
[…]
[inst2]organization=NASA Ames Research Center, Earth Science Division,,addressline=Moffett Field, state=California, postcode=94035, country=USA
1 Introduction
As climate change accelerates, forests represent a key ecosystem to protect as they act as one of the main terrestrial carbon sinks (Keenan and Williams, 2018), a shelter for a major part of Earth’s biodiversity (Lindenmayer and Franklin, 2013) , and a critical environment for numerous fragile human communities (Fernández-Llamazares et al., 2021). In particular, the boreal ecosystem is a subarctic biome in the high northern latitudes characterized by coniferous and mixed deciduous-coniferous forests. They represent one of the largest terrestrial carbon sinks, with approximately $367.3$ petagrams to $1715.8$ petagrams of carbon stored (Bradshaw and Warkentin, 2015). However, they are at risk of permafrost thaw due to land impacts (Li et al., 2021) and are increasingly subject to long and devastating wildfire seasons (McCarty et al., 2021).
While wildfires severity is amplifying globally (areas burned by forest fires have seen a steady yearly increase of $\sim 5\%$ since 2001 (Tyukavina et al., 2022)), its effect is particularly devastating for the boreal ecosystem, representing roughly $70\%$ of the fire-related tree cover loss (Tyukavina et al., 2022) and where single wildfire events, like those during the 2023 Canadian wildfires season, can compete with annual $\text{CO}_{2}$ emissions of major industrialized nations (Byrne et al., 2024). Locally, boreal wildfires have a direct impact on the land surface as they directly increase permafrost thaw (Li et al., 2021; Zhao et al., 2024), and contribute to vegetation shifts to more fire-prone grassland-/steppe-dominant landscapes, as well as dry peat (Zhao et al., 2024; McCarty et al., 2021). Boreal forests span $58\%$ of Canada’s land mass: in this paper we use Canada as the area of interest to tackle the problem of wildfire forecasting in boreal ecosystems.
The use of remote sensing data to map and monitor wildfires has expanded, with studies considering satellite-based observations of vegetative fuel conditions, individual fire events, and the impacts of smoke. Numerous wildfire tools exist, with three main use cases focused on the different phases of a wildfire: before a wildfire occurs (pre-fire), during a wildfire (active fire), and after (post-fire). For active fires, satellite products and models detect active fire ”hotspots” in near real-time (de Almeida Pereira et al., 2021; Růžička et al., 2022) or predict wildfire spread (Huot et al., 2022; Hoang et al., 2022), providing tools for wildfire management and decision-making. In the aftermath of wildfires (post-fire), methods to precisely segment the perimeter of burned areas (Hu et al., 2023; Zhang et al., 2022) were developed to evaluate wildfires emissions, from $\text{CO}_{2}$ to black carbon, and estimate the impact of wildfires on the local ecosystem, natural resources, and communities. This paper focuses on the pre-fire phase, shaped from a methodological perspective as a forecasting task. Wildfire forecasting, often referred to as wildfire susceptibility or likelihood modeling (Pelletier et al., 2023; Zhang et al., 2021), aims to predict the spatial probability of a wildfire occurring in a given time horizon and at a given spatial resolution. This is done by producing wildfire probability maps. Wildfire forecasting is particularly useful for wildfire management by supporting staff and resource planning (Wotton, 2009).
Wildfire forecasting is, by definition, a difficult task since it seeks to represent a complex and stochastic phenomenon. Fire susceptibility depends on several drivers: i) hydrometeorological conditions are the main variables that impact the suitability of vegetative fuels for combustion (ie, ’dryness’) and fire propagation (Krawchuk et al., 2009). Fire susceptibility is also directly linked to ii) the available biomass for combustion, which depends on iii) the type of vegetation and other indicators such as iv) dead or live fuel moisture (Krawchuk et al., 2009). Climate change makes those predictors for wildfire occurrence non-stationary. For example, meteorological patterns are highly variable in the boreal biome, which can lead to extreme fire seasons (McCarty et al., 2021). In addition, numerous vegetation changes have been observed or predicted, such as permafrost thaw (Li et al., 2021; Zhao et al., 2024), peatland destruction (Bourgeau-Chavez et al., 2022), or a shift from coniferous to deciduous forest (McCarty et al., 2021). Moreover, for effective wildfire forecasting, it is necessary to estimate v) the probability of ignition caused by humans or lightning (Pérez-Invernón et al., 2023) through proxies such as the proximity to human settlements. This variability implies that for similar environmental conditions, a wildfire may or may not occur based on latent variables for the model, which makes wildfire forecasting an especially complex task where ignition, even from lightning-only, is hard to model (Coughlan et al., 2021; Bates et al., 2021).
Historically, wildfire forecasting was performed by producing fire weather indices, such as the Canadian Forest Fire Weather Index (FWI) or the National Fire Danger Rating System (NFDRS), which are mainly driven by meteorological and fuel moisture data. Fire weather indices aim to represent complex relationships between fire predictors through constrained equations based on simplifying assumptions, such as the forest type (e.g ”Pinus Banksiana”) and neglecting the topography, leading to necessary recalibrations of the indices for specific areas (Steinfeld et al., 2022; De Jong et al., 2016). For instance, across Canada, the FWI has limitations in properly identifying the hydrometeorological conditions for combustion across all land cover types, and particularly so in peatlands (Waddington et al., 2012). Moreover, those indices cannot approximate the stochastic character of wildfire occurrence as they focus on flammability conditions. In parallel, traditional machine learning (ML) algorithms based on handcrafted features were proposed in (Martell et al., 1987, 1989) to identify drivers linked to wildfire occurrence, such as hydrometeorological conditions and human activities. These ML algorithms are limited in their ability to represent complex predictor relationships and possible spatio-temporal patterns, mostly because of the rigidity of the features used. Nevertheless, these methods are still widely utilized (Buch et al., 2023; Rodrigues et al., 2022) even in remote sensing (Maffei et al., 2021; Chowdhury and Hassan, 2015).
The growing availability of open-access remote sensing data (Reichstein et al., 2019; Camps-Valls et al., 2021), which enables the monitoring of large and remote regions, now allows mapping the drivers of fire susceptibility. This accumulation of data contributed to the emergence of wildfire forecasting models leveraging remote sensing imagery with neural networks (Xu et al., 2025). When processing hydrometeorological data in the form of one-dimensional inputs (i.e. tabular data), the method of choice is the Multi-Layer Perceptron (MLP) (Buch et al., 2023; Bakke et al., 2023; Milanović et al., 2020), while for temporal series of tabular inputs, Long Short Term Memory (LSTM) networks have been proposed (Natekar et al., 2021) due to their ability to capture temporal relationships. When considering spatial inputs, convolutional neural networks (CNN) and vision transformer (ViT) have been explored (Prapas et al., 2022, 2023). Finally, for spatio-temporal data, architectures like convolutional LSTM (ConvLSTM) (Kondylatos et al., 2022; Huot et al., 2020; Prapas et al., 2021; Bali et al., 2021), which join the sequence processing abilities of LSTM networks to the spatial awareness of CNNs, have been proposed. There is no clear consensus on the best model to use, as results seem to vary depending on the dataset characteristics (Kondylatos et al., 2022; Huot et al., 2020; Prapas et al., 2021; Jain et al., 2020), region of interest, forecast horizon, and predictors list. In terms of data, most methods leverage hydrometeorological predictors, from reanalysis data like ERA5 or weather stations with spatial resolution varying from $\sim 27$ km to $4$ km. Finally, researchers resort to remote sensing products (characterizing the vegetation) at higher resolution (up to $500$ m) and static factors symptomatic of land cover type and human activities (Kondylatos et al., 2022; Prapas et al., 2021; Bakke et al., 2023; Prapas et al., 2022; Bali et al., 2021). The individual predictors are then re-sampled to the target resolution corresponding to the final wildfire probability map, varying from 0.250.25° for global applications (Bakke et al., 2023; Prapas et al., 2022) to up to $1$ km for localized regions (Kondylatos et al., 2022; Huot et al., 2020; Prapas et al., 2021). For instance, in the context of large countries such as Canada, which spans thousands of kilometers, the output resolution of current wildfire probability maps is $\sim 0.1$° (Bali et al., 2021). This represents an important limitation, as such coarse wildfire probability maps prevent wildfire management from properly allocating resources at a finer scale and lead to the underestimation of potential smaller wildfires.
In this paper, we propose a multi-modal and spatio-temporal dataset covering Canada to enable high-resolution ($100$ m) wildfire forecasting and benchmark different models to demonstrate their potential. Our contributions are as follows:
  1. 1.
    We introduce a benchmark dataset, CanadaFireSat [^1] Code for the data generation and the model benchmarking can be accessed, respectively, at github.com/eceo-epfl/CanadaFireSat-Data and github.com/eceo-epfl/CanadaFireSat-Model, available on the HuggingFace Hub, for high-resolution wildfire forecasting at $100$ m over Canada in 8-day forecasting window. CanadaFireSat enables high-resolution wildfire forecasting by resorting to temporal series of multi-spectral images (Sentinel-2) complemented by temporal series of environmental drivers from both reanalysis data (ERA5) and coarse resolution satellite products (MODIS), as shown in Figure 1.
  2. 2.
    We investigate the impact of negative sampling on wildfire forecasting through the collection of two test sets across the 2023 extreme wildfire season for CanadaFireSat. Besides a classic test set following the same sampling strategy as the train and validation sets, where wildfire forecasting models show compelling performance, we also propose a hard test set sampled adversarially: this allows studying the lower-bound performance of models under extreme conditions, where ignition constitutes the key discriminating factor to identify potential wildfires.
  3. 3.
    We demonstrate the potential of learning multi-modal models for high-resolution wildfire forecasting by benchmarking two state-of-the-art computer vision architectures on CanadaFireSat: ResNet (He et al., 2016) and ViT (Dosovitskiy, 2020) across three settings with varying input modalities: [Uncaptioned image]satellite images only (Sentinel-2 at $10$ m), [Uncaptioned image]environmental predictors only (ERA5 at $11$ km, FWI at 0.250.25°, MODIS at $1$ km and $500$ m), and [Uncaptioned image]satellite and environmental data.
CanadaFireSat allows a big leap in terms of resolution with respect to what was possible with previous datasets, such as (Huot et al., 2020) or (Prapas et al., 2021), both using a target resolution of $1$ km over the U.S. and the Eastern Mediterranean region, respectively. Moreover, our results on CanadaFireSat demonstrate that: i) deep learning models outperform a knowledge-driven baseline (FWI) in both normal and extreme fire seasons, and ii) multi-spectral and hydrometeorological data complement each other, with multi-modal models providing the most accurate predictions.
Figure 1: The CanadaFireSat benchmark and the high-resolution wildfire forecasting task.
Statistic Value
Total Samples 177,801
Target Spatial Resolution 100 m
Region Coverage Canada
Temporal Coverage 2016 - 2023
Sample Area Size $2.64\;\text{km}\times 2.64\;\text{km}$
Fire Occurrence Rate 39% of samples
Total Fire Patches 16% of patches
Training Set (2016-2021) 78,030 samples
Validation Set (2022) 14,329 samples
Test Set (2023) 85,442 samples
Sentinel-2 Temporal Median Coverage 55 days (8 images)
Number of Environmental Predictors 58
Data Sources ERA5, MODIS, CEMS
Table 1: Main Statistics of the CanadaFireSat Dataset
2 The CanadaFireSat Dataset
In this section, we present CanadaFireSat, a benchmark dataset for high-resolution wildfire forecasting. First, we describe the sampling scheme for the selection of positive and negative data samples in Section 2.1. Then, in Section 2.2 we detail the set of predictors extracted and combined to build our multi-modal learning benchmark for high-resolution wildfire forecasting. Table 1 summarizes CanadaFireSat’s main characteristics.
2.1 Sample Identification
Covering the entirety of Canada with Sentinel-2 images at $10$ m requires extremely high storage capacity, beyond the size of typical datasets. As such, to represent all territories and provinces of Canada, we build CanadaFireSat by resorting to a sampling strategy. As our fire labels are binary, we sample the dataset as a series of positive and negative examples. For fire (positive) sample identification, we first extract all fires that occurred between 2015 and 2023, as described in Section 2.1.1. Then, samples not including any fire event (negative) are sampled across the same period for all provinces and territories, depending on their FWI and acquisition dates, as detailed in Section 2.1.2.
Figure 2: Burned area in Canada in millions of hectares extracted from NBAC, compared to the values reported by the Canadian Interagency Forest Fire Centre (CIFFC). (a) shows the annual burned area for Canada from 2016 to 2023. The difference between CIFFC and NBAC reported burned area has multiple explanations. First, the CIFFC statistics are not standardized across all territorial fire management agencies, contrary to NBAC. This is directly linked to data collection timelines, as CIFFC may provide near-real-time estimates while NBAC is compiled up to 6 months after the calendar year, leaving more room for comprehensive post-fire analysis. (b) reports the per-region burned area for 2023 only, where the most impacted provinces and territories were Québec, Northwest Territories (Natural Resources Canada, which provides NBAC data, includes Nunavut fires in Northwest Territories statistics), and British Columbia. We note that the most impacted regions are those with the strongest discrepancies between reported numbers from CIFFC and NBAC.
2.1.1 Positive Samples
Fire samples in our CanadaFireSat dataset are identified based on the fire polygons of the National Burned Area Composite [^2] Available at https://cwfis.cfs.nrcan.gc.ca/datamart/metadata/nbac (NBAC) (Hall et al., 2020) from the Canadian National Fire Database. NBAC has been compiled annually since 1972 and integrates data from Natural Resources Canada, provincial and territorial agencies, and Parks Canada, using a rule-based approach to select the most accurate data source to delineate the burned area perimeters; this includes ground and aerial surveys or post-event satellite imagery analysis from Landsat (5, 6, 7, 8, 9 or MSS), Sentinel-2, MODIS, VIIRS, and AVHRR. We focus on all fires since 2015 (the launch of the first Sentinel-2 satellite) up to 2023, with no restriction on ignition sources or other fire metadata. Over this time, a large majority of the polygons were compiled from ground survey, Landsat, aerial survey, and Sentinel-2 in this respective order. In Figure 2a, we report the NBAC yearly average burned area for this period, with 2022 reaching $1.38$ mha burned and 2023 reaching $14.01$ mha burned. This outlines the difference in wildfire season severity for our validation (2022) and test (2023) sets compared to the average from 1972 to 2015 of $\sim 2.03$ mha. In other words, 2023 was an exceptional fire season.
Positive samples for the CanadaFireSat dataset are extracted from the NBAC fire polygons through two aggregation processes. First, through a spatial aggregation on a $2.8\;\text{km}\times 2.8\;\text{km}$ grid over Canada, where positive samples are identified as the grid entries intersecting the fire polygons. We used a small buffer around the $2.64\;\text{km}\times 2.64\;\text{km}$ Sentinel-2 tiles to avoid any potential overlap between samples due to imprecision in the data processing. Second, a temporal aggregation is performed in two steps: 1) all fires temporally overlapping inside a grid entry are accounted as a single fire occurring from the first fire start date to the last fire end date, and 2) leveraging the 8-day temporal grid from products such as NDVI from MODIS (starting each year at the 1st of January) we aggregate all fires within a spatial grid entry occurring during the same 8-day window. This is done to build our 8-day wildfire forecasting benchmark where, for a given time-step $t$, our model should predict the probability of a fire occurring in the next 8 days, i.e. from $t$ to $t+7$ included, leveraging predictors (both satellite and environmental, see Section 2.2) from $t-\Delta t$ to t1t-1. In Figure 3(a), we showcase the spatial distribution of positive samples across Canada, for a total of $n_{pos}=88,110$ samples before any post-processing (detailed in Section 2.2). Outside of British Columbia, most fires occur in the boreal ecosystem. This pattern is very visible across Alberta, Saskatchewan, and Manitoba, where the Great Plains in the southern portions of these provinces show little to no fires compared to the boreal forest in the north.
(a) Positive sample distribution across the period 2015-2023
(b) Negative sample distribution across the period 2015-2023
Figure 3: Distribution of positive (containing burned area) and negative samples (following our FWI-based sampling strategy) from 2015-2023, before any post-processing.
2.1.2 Negative Samples
As we aim to build our benchmark on multi-modal inputs, including satellite image time series, we are limited in disk storage to densely sample Canada over the whole period from 2015 to 2023. Therefore, we sample a negative set of size $n_{neg}=2\cdot n_{pos}$ to match the degree of imbalance of other wildfire forecasting datasets (Huot et al., 2020; Prapas et al., 2021; Kondylatos et al., 2022). We sample from the same grid defined in Section 2.1.1, $G_{y,r}$, for each year $y$ between the first and last fires during that year (so beyond the wildfire season), and across all regions $r$. For a given year $y>2015$ and region $r$, we avoid locations where a fire occurred in the previous years: $\bigcup_{i=2015}^{y-1}F_{i,r}$, or locations that were already selected as negative samples in the previous years: $\bigcup_{i=2015}^{y-1}N_{i,r}$. Our negative set for a given region and year can be defined as:
$N_{y,r}\sim S_{y,r}=\{x\in G_{y,r}|\;x\notin\bigcup_{i=2015}^{y-1}F_{i,r}\wedge x\notin\bigcup_{i=2015}^{y-1}N_{i,r}\}$ (1)
where $N_{y,r}$ is the set of negative samples and $S_{y,r}$ the set of potential locations in the grid. We sample $N_{y,r}$ uniformly across levels (defined by decile bins) of the FWI:
$P_{FWI}(x|N_{y,r})\propto P_{FWI}(x|S_{y,r})$ (2)
In practice, this is done by partitioning the FWI distribution into ten decile bins: $B_{1},\dots,B_{10}]$ across the FWI quantiles $Q_{1},\dots,Q_{9}]$ such that each bin contains approximately $10\%$ of the observations, and uniformly sampling across those decile bins for $N_{y,r}$. Each bin $B_{l}$ is defined as a subset of the FWI range:
$B_{l}=\left\{x\in\text{FWI}\,\middle|\,Q_{l-1}<x\leq Q_{l}\right\},\quad\text{for}\>l=1,\dots,10$ (3)
where $Q_{0}=0$, and $Q_{10}=+\text{inf}$ are the bounds of the FWI range. This way, the negative population is representative of all fire weather conditions for each region and year, including cases where a high FWI was predicted, but no fire was observed.
Figure 3(b) presents the spatial distribution of the sampled negative locations across all years: it shows that, per region, the negative samples are well spread spatially, contrary to the positive samples, as we aim to represent the complete patterns of fire danger conditions. British Columbia, Alberta, Saskatchewan, and Manitoba contain the highest concentration of negative samples in certain areas due to the high concentration of fires in those regions (negative samples are sampled twice as much as positive ones). On the contrary, Nunavut, Newfoundland and Labrador, and New Brunswick are less densely sampled due to a lack of fires during the analyzed period. We select in total $n_{neg}=176,650$ negative samples that, combined with our positive samples $n_{pos}$, consitute the CanadaFireSat Train (2016 - 2022), Val (2022), and Test (2023) sets. Note that some of these samples will be filtered out through the post-processing procedure described in Section 2.2.
In Figure 4(a), we present the annual FWI mean for the negative sample set. We see that up to the decile bin number $4$ with a FWI mean: $\overline{\text{FWI}}=0.62$, most negative samples will have an FWI close to 0, as the FWI distribution of available locations for negative samples consistently presents an important peak in this range. This is representative of the FWI conditions across all regions of Canada between the first and last fires of each year. Furthermore, we show in Figure 4(b) and 4(c) that across two commonly impacted regions by wildfires (Alberta and the Northwest Territories), there are strong differences in FWI decile bin mean value between regions, with the delta for the top decile bin reaching up to $\sim 14$ in 2021. This can be explained by the higher latitude of the Northwest Territories compared to Alberta and the presence of permafrost in their northernmost areas.
(a) Canada
(b) Alberta
(c) Northwest Territories
Figure 4: Annual FWI mean over four decile bins: $\{1,4,7,10\}$ across Canada, Alberta, and Northwest Territories.
We also observe a strong inter-annual variability between 2022 and 2023, as the latter was a record-breaking wildfire season in Canada (Jain et al., 2024), resulting in $19.6\%$ of fire patches in the Test set compared to $11.5\%$ in the Val set. This distribution shift shows that, despite similar fire weather conditions as presented in Figure 4(a), wildfire occurrence is significantly higher in the Test set compared to the Training set. This can lead to the overestimation of the performance of wildfire forecasting models: by looking at the distribution of positive and negative samples in Figure 5, one can observe that the FWI alone is a highly discriminative feature for the class fire (see Section 4.1). As a result, we introduce an adversarial sampling strategy for the negative samples to study the lower-bound performance of wildfire forecasting models for the extreme year $2023$, named Test Hard. In this adversarial test set, we aim to make the distribution of the negative population similar to that of the positive population with respect to the FWI, making ignition the main discriminative factor. To sample negative samples for Test Hard, we perform a stratified sampling for the year 2023 in the following way. First, by extending Equation 1 to account for both land cover and the month of the year. Then, for a given land cover $c$ and month of the year $m$, we sample $N_{y,r,m,c}$ uniformly across levels (defined by decile bins) of the FWI for the positive samples population $F_{y,r,m,c}$ :
(a) FWI distribution in log-scale for the Test set across positive and negative samples.
(b) FWI distribution in log-scale for the Test Hard set across positive and negative samples.
Figure 5: Comparison of the FWI distribution in log-scale across the Test and Test Hard sets for both positive and negative samples.
$P_{FWI}(x|N_{y,r,m,c})\propto P_{FWI}(x|F_{y,r,m,c})$ (4)
and sample $n_{neg}(y,r,m,c)\simeq 2\times n_{pos}(y,r,m,c)$ negatives. The land cover is downloaded from ESA WorldCover at $10$ m for 2020. The resulting distribution is shown in Figure 5 and represents 77,247 complementary negative samples to CanadaFireSat statistics reported in Table 1. By deploying the trained networks on Test Hard, where ignition acts as the main distriminative factor, we can assess their performance on modeling this complex triggering factor whose patterns can only be implicitly learned from the training data. For this reason, the performance of our trained models for high-resolution wildfire forecasting on Test Hard can be considered as a lower bound for such an extreme wildfire season as presented in Section 4.1. Further details about the distribution of samples across land-cover classes are provided in Figure 17.
2.2 Predictors
The predictors used in CanadaFireSat fall into two categories: satellite image time series and environmental data.
2.2.1 Satellite Image Time Series
S-2 Input
Refer to caption Refer to caption Refer to caption Refer to caption
S-2 Input
Refer to caption Refer to caption Refer to caption Refer to caption
S-2 Input
Refer to caption Refer to caption Refer to caption Refer to caption
S-2 Post-Fire
Refer to caption Refer to caption Refer to caption Refer to caption
S-2 and Label
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6: Row 1-3 Samples of Sentinel-2 input time series for four locations in Canada. We show only the RGB bands at $10$ m resolution with rescaled intensity. Row 4 Sentinel-2 images after the fire occurred. Row 5 Fire polygons used as labels with the Sentinel-2 images post-fire as background.
To be able to forecast wildfires at a patch resolution of $100$ m, we need high-resolution information. However, hydrometeorological fire danger predictors cannot be found at $100$ m resolution for the entirety of Canada. Therefore, we investigate the potential of multi-spectral high-resolution satellite images as proxies for fire predictors following previous literature (Pelletier et al., 2023; Yang et al., 2021). We use the 13 bands from Sentinel-2 (S2) L1C harmonized data as proxies to several known fire predictors, such as NDVI or soil moisture. We use the L1C products as they are directly available for the whole period 2015-2023 without any need for further processing. Moreover, we extract temporal data to better estimate the impact of changes in the hydrometeorological conditions on the local ecosystem. Top-of-atmosphere reflectance from Sentinel-2 is impacted by aerosols, clouds, topography effects, and other phenomena that can bias its measurement across the multi-spectral bands for numerous land cover types (Sola et al., 2018), in particular for shorter wavelengths like the RGB bands. This can impact the computation of radiometric indices often used in burned area mapping (Howe et al., 2022) and the precision of non-local and multi-temporal analyses of Sentinel-2 data. However, machine and deep learning can largely mitigate those limitations by implicitly learning the approximate corrections necessary for the downstream application targeted through correction agnostic models (Rußwurm et al., 2023; Wright et al., 2025) or L1C specific ones (Medina-Lopez, 2020; Wright et al., 2024) even in the context of burned area mapping (Rumora et al., 2020). We hypothesize that our models can mitigate the lack of atmospheric corrections for the task of wildfire forecasting.
For a given sample $x_{t}\in G_{y,r}$ (the $2.8\;\text{km}\times 2.8\;\text{km}$ grid over Canada), we download all the full (no missing values) S2 images of size $2.64\;\text{km}\times 2.64\;\text{km}$ following (Manas et al., 2021) centered within each grid cell $x_{t}$ between the date t64t-64 days and t1t-1 day. We exclude images with a cloud cover above $40\%$. This represents 13 images, given the average revisit time of 5 days for S2 (after the launch of Sentinel-2B). To avoid artifacts, we use a lossless compression, and we multiplied each band intensity by a factor $1e^{-4}$ to then rescale the values to 8-bit unsigned integers. Once all S2 images are extracted, a second filter on cloud coverage was applied, based on the S2 cloud probability product, but focusing only on the sample location. After this filtering, the samples $x_{t}$ with less than three S2 images or covering a period of less than 40 days are removed, as we aim to learn local temporal dynamics. Finally, as multiple S2 tiles can cover a sample $x_{t}$, we keep the tile with the most valid images for this sample. The final positive sample set after filtering is of size $69,876$ ($\sim 79\%$ of the original set), and the negative sample set is of size $107,925$ ($\sim 61\%$ of the original set). For Test Hard, the final number of negative samples is $66,406$ ($\sim 86\%$ of the previously identified samples).
The fire polygons associated with each positive sample are rasterized based on the B3 band of the S2 image that preceded the start of the fire. This process outputs binary maps of size $264\times 264$ pixels at a resolution of $10$ m that will then be downscaled to $100$ m resolution during training. Figure 6 shows examples of S2 image time series from the positive set and the expected output when a fire occurred (last row).
2.2.2 Environmental Predictors
Fire weather indices and most wildfire forecast models rely on hydrometeorological drivers such as temperature, precipitation, and soil moisture. Some forecast models also leverage vegetation indices such as NDVI, EVI, or LAI. We include such coarse environmental predictors (summarized in Table 2) despite the difference in resolution between them and our target outputs, since we believe that multi-modal methods can benefit from them, as they are strongly correlated to fire probability.
  • 1.
    First, we extract five different MODIS products from MOD15A2H, MOD11A1, and MOD13A1 that describe the vegetation state and temperature at moderate-to-coarse-resolution: $500$ m and $1$ km. Vegetation indices are 8-day or 16-day composites, which, similarly to SeasFire (Prapas et al., 2022), drive the temporal aggregation over 8 days of the other environmental predictors, and the NBAC burned area polygons.
  • 2.
    We also extracted 12 hydrometeorological drivers from ERA5-Land daily (detailed in Table 2) at coarse-resolution ($11$ km), and aggregated those variables through mean, max, and min operators on the 8-day temporal grid defined by MODIS. We extend this set of predictors with three additional ones: relative humidity, vapor pressure deficit, and wind speed, computed locally from ERA5 data.
  • 3.
    Lastly, we leverage indices related to fire danger from the Copernicus Emergency Management Service (CEMS): FWI, also used in negative sampling, and drought code, both from the Canadian Forest Fire Weather Index. This data is the coarsest of all our environmental predictors with a resolution of 0.250.25° for both latitude and longitude.
These predictors are then post-processed to set to NaN any extreme values and aligned both spatially and non-spatially with our positive and negative samples. Similar to the satellite image time series, for each sample $x_{t}$, we extract the environmental predictors from t64t-64 days to t1t-1. The non-spatial alignment is done via the weighted average of a given predictor over the target grid cell. The spatial alignment is done for each predictor by extracting a small window of data centered on $x_{t}$. The window size varies depending on the source resolution. We extract windows of dimension $(32,32)$ for MODIS products at $500$ m and $(16,16)$ for MODIS product at $1$ km. Moreover, for ERA5-Land data, we extract windows of size $(32,32)$, and $(13,13)$ for CEMS. As a consequence, for a given sample $x_{t}$ CanadaFireSat provides spatial predictors at multiple scales covering different spatial contexts. Models trained on CanadaFireSat should consider this difference in scale across modalities, as those presented in Section 3.
Dataset Name Units Aggregation Resolution Source MODIS NDVI - 16-day composite 500 m Google Earth Engine EVI - 16-day composite 500 m Google Earth Engine LST Day (1km) K 8-day mean, max, min 1 km Google Earth Engine FPAR - 8-day composite 500 m Google Earth Engine LAI - 8-day composite 500 m Google Earth Engine ERA5-Land Surface Pressure Pa 8-day mean, max, min 11.1 km Google Earth Engine Total Precipitation Sum m 8-day mean, max, min 11.1 km Google Earth Engine Skin Temperature K 8-day mean, max, min 11.1 km Google Earth Engine U Component of Wind (10m) m/s 8-day mean, max, min 11.1 km Google Earth Engine V Component of Wind (10m) m/s 8-day mean, max, min 11.1 km Google Earth Engine Temperature (2m) K 8-day mean, max, min 11.1 km Google Earth Engine Temperature (2m, Max) K 8-day mean, max, min 11.1 km Google Earth Engine Temperature (2m, Min) K 8-day mean, max, min 11.1 km Google Earth Engine Surface Net Solar Radiation Sum J/m² 8-day mean, max, min 11.1 km Google Earth Engine Surface Solar Radiation Downwards Sum J/m² 8-day mean, max, min 11.1 km Google Earth Engine Volumetric Soil Water Layer 1 m³/m³ 8-day mean, max, min 11.1 km Google Earth Engine Dewpoint Temperature (2m) K 8-day mean, max, min 11.1 km Google Earth Engine Relative Humidity % 8-day mean, max, min 11.1 km Own Calculation Vapor Pressure Deficit hPa 8-day mean, max, min 11.1 km Own Calculation Wind Speed (10m) m/s 8-day mean, max, min 11.1 km Own Calculation CEMS Drought Code - 8-day mean, max, min 0.25° ( 28 km) CEMS Early Warning Data Store Fire Weather Index - 8-day mean, max, min 0.25° ( 28 km) CEMS Early Warning Data Store
Table 2: Overview of the environmental predictors.
3 Methods
To demonstrate the feasibility of forecasting wildfires at $100$ m resolution, we benchmark two deep learning architectures on the proposed CanadaFireSat dataset. We chose a CNN and a Transformer as representative computer vision models, whose encodings are used to forecast wildfire probability at an 8-day horizon. To account for multi-modal interactions, models are trained in three different settings: [Uncaptioned image]satellite images only (Sentinel-2), [Uncaptioned image]environmental predictors only (ERA5, CEMS, MODIS), and when both [Uncaptioned image]satellite and environmental data are available. Detailed information on the settings can be found in Table 3.
Setting Source Format Type
[Uncaptioned image] Sentinel-2 Spatial Multi-Spectral Images
[Uncaptioned image] MODIS Spatial Environmental Products
ERA5-Land Spatial Climate Reanalysis
CEMS Spatial Fire Indices
[Uncaptioned image] Sentinel-2 Spatial Multi-Spectral Images
MODIS Tabular Environmental Products
ERA5-Land Tabular Climate Reanalysis
CEMS Tabular Fire Indices
Table 3: Descriptions of the modality settings for the training of the wildfire forecasting models.
For CanadaFireSat, wildfire forecasting is framed as a binary classification task (fire vs no fire) at the patch level, i.e., a binary patch classification. Across our experiments, the original labels at a native resolution of $10\;\text{m}\times 10\;\text{m}$ are re-scaled to $100\;\text{m}\times 100\;\text{m}$, by labeling a patch with the binary class fire if any pixel within the patch is labeled as burned. This design decision aims to focus on providing alerts for any size of fires at the expense of false positive pixels at the native resolution and is often used in wildfire prediction at both coarse (Prapas et al., 2022; Bakke et al., 2023) and high-resolution (Pelletier et al., 2023). It is also motivated by the shortcomings of MODIS, in particular the MCD64A1 burned area product, which is recurrently used in coarse wildfire forecasting (Huot et al., 2020; Rodrigues et al., 2022; Prapas et al., 2021) despite underestimating burned area (Bakke et al., 2023; Zhu et al., 2017).
Finally, as satellite image time series are not evenly spaced due to cloud cover, we add as complementary information the day of the year for all our predictors composing the time series. Other details on the experimental setup for all architectures can be found in B. We analyze the impact of satellite image time series on model performance in C.
3.1 CNN-based Architecture
In the CNN-based architecture, satellite image time series are processed in a factorized manner: first spatially and then temporally, for both settings [Uncaptioned image]satellite images only and [Uncaptioned image]satellite and environmental data, as shown in Figure 12 and Figure 7, respectively. For a given satellite image time series $x_{1:T}=\{x_{t}\}_{t=1}^{T}$, with $x_{t}\in\mathbb{R}^{H\times W\times C}$ being a single time step with $C$ the number bands and the day of the year, and $T$ a fixed number of time steps, each image $x_{t}$ is first encoded independently by a ResNet-50 pre-trained on ImageNet (He et al., 2016): $f(x_{t})=\{z_{i,t}\}_{i=1}^{N_{S}}$, with, $z_{i,t}\in\mathbb{R}^{H_{i}\times W_{i}\times D_{i}}$, which outputs $N_{S}=3$ feature maps of channel dimension $D_{i}$, each feature map corresponding to a different scale. The encoding of all time steps is done in parallel, and each scale-specific feature map, $z_{i,t}$, is concatenated independently for each scale across the temporal axis: $z_{i,1:T}=\{z_{i,t}\}_{t=1}^{T}$. Then, the spatio-temporal encoding is done via one ConvLSTM model per scale. By extracting the last hidden state from each ConvLSTM: $g_{i}$, we obtain feature maps $g_{i}(z_{i,1:T})=s_{i}$, with $s_{i}\in\mathbb{R}^{H_{i}\times W_{i}\times D^{\prime}_{i}}$ at 3 different scales with channel dimension $D^{\prime}_{i}<D_{i}$, providing multiple levels of contextual information.
In setting [Uncaptioned image]satellite images only (Figure 12), our final multi-scale feature maps $\{s_{i}\}_{i=1}^{N_{S}}$ are passed to a U-Net-like decoder. The output of the decoder is interpolated to the dimensions of the label feature map: $H_{\text{fire}}=W_{\text{fire}}=\frac{H}{10}=\frac{W}{10}$ to match the patch resolution. This is finally passed to binary patch classification layer to output the class probabilities: $h({\{s_{i}\}_{i=1}^{N_{S}}})=\hat{y}\in[0,1]^{H_{\text{fire}}\times W_{\text{fire}}\times 2}$, with $h$ the function representing the decoder, interpolation, and patch classification layer.
In the multi-modal setting [Uncaptioned image]satellite and environmental data (Figure 7), the above architecture is extended to process in parallel the time series of non-spatial environmental predictors $x_{\text{env},1:T_{\text{env}}}=\{x_{\text{env},t}\}_{t=1}^{T_{\text{env}}}$, with $x_{\text{env},t}\in\mathbb{R}^{N_{\text{env}}}$ being the data for a single time step with $N_{\text{env}}$ environmental predictors. $T_{\text{env}}$ is a fixed number of time steps. Following (Gorishniy et al., 2022) for tabular data encoding, each environmental predictor is projected to a high-dimensional space with specific MLP layers: $\forall j\in\{1,\dots,N_{\text{env}}\},\;f_{\text{env},j}(x_{\text{env},t}^{j})=z_{\text{env},t}^{j}$, with $z_{\text{env},t}^{j}\in\mathbb{R}^{D_{\text{env}}}$. The projected features are then averaged across the $N_{\text{env}}$ dimension to obtain $z_{\text{env},1:T_{\text{env}}}\in\mathbb{R}^{D_{\text{env}}\times T_{\text{env}}}$ and passed to an LSTM model for temporal encoding $g_{\text{env}}(z_{\text{env},1:T_{\text{env}}})=s_{\text{env}}\in\mathbb{R}^{D_{\text{env}}}$ , which we use as the final environmental encoded feature. This one-dimensional vector is replicated spatially and concatenated with the final feature map from the U-Net-like decoder before the patch classification layer: $h({\{s_{i}\}_{i=1}^{N_{S}}},s_{\text{env}})=\hat{y}\in[0,1]^{H_{fire}\times W_{fire}\times 2}$.
Figure 7: CNN Architecture for Wildfire Prediction in setting Refer to captionsatellite and environmental data. Top: Satellite image time series encoding, Bottom: Environmental predictors encoding.
For setting [Uncaptioned image]environmental predictors only (Figure 13), where the model is trained using only environmental predictors at a resolution varying from $500$ m up to $28$ km (see Table 2), we first split the predictors into two groups: mid-resolution inputs $x_{\text{mid},1:T_{\text{env}}}=\{x_{\text{mid},t}\}_{t=1}^{T_{\text{env}}}$, with single time step $x_{\text{mid},t}\in\mathbb{R}^{H_{\text{mid}}\times W_{\text{mid}}\times N_{\text{mid}}}$ for all MODIS data, and low-resolution inputs for all ERA5 and CEMS data: $x_{\text{low},1:T_{\text{env}}}=\{x_{\text{low},t}\}_{t=1}^{T_{\text{env}}}$, with single time step $x_{\text{low},t}\in\mathbb{R}^{H_{\text{low}}\times W_{\text{low}}\times N_{\text{low}}}$. In this setting, we leverage spatial environmental inputs and not tabular to compensate for the absence of high-resolution satellite imagery from Sentinel-2, which provides spatial context for settings [Uncaptioned image]satellite images only and [Uncaptioned image]satellite and environmental data. In each group, as not all predictors have the same spatial resolution (see Table 2), we upsampled all predictors to the highest available resolution. Details on the different spatial dimensions for each group can be found in the B. We partially modify the architecture from setting [Uncaptioned image]satellite and environmental data, as shown in Figure 13. First, mid-resolution inputs are used as an alternative to satellite image time series. In practice, all the satellite image processing model components stay the same for the mid-resolution inputs group: the spatial encoding $f$, the scale-specific temporal encoding $g_{i}$, and the final head $h$ (corresponding to the decoder, interpolation layer, and patch classification layer). We simply extend the number of multi-scale feature maps to $N_{S}=5$ because of the lower resolution of the input data. Moreover, we exchange ConvLSTM with LSTM when the output feature maps from $f$ become one-dimensional (for $i=5$). The second branch of the model is adapted to process spatial data for the low-resolution inputs group. We use a smaller pre-trained CNN architecture to encode independently each time step, similarly to the processing of satellite images described above: we use ResNet-18 (He et al., 2016) to obtain a one-dimensional feature vector $f_{\text{low}}(x_{\text{low},t})=z_{\text{low},t}$, with $z_{\text{low},t}\in\mathbb{R}^{D_{\text{low}}}$. The temporally concatenated features $z_{\text{low},1:T_{\text{env}}}\in\mathbb{R}^{D_{\text{low}}\times T_{\text{env}}}$ are passed to an LSTM model $g_{\text{low}}(z_{\text{low},1:T_{\text{env}}})=s_{\text{low}}$, with $s_{\text{low}}\in\mathbb{R}^{D^{\prime}_{\text{low}}}$ to obtain low-resolution encoded features with $D^{\prime}_{\text{low}}<D_{\text{low}}$. Similarly to the multi-modal architecture, this one-dimensional vector is replicated spatially and concatenated with the final feature map from the U-Net-like decoder that has processed the mid-resolution group.
In all three settings, the training is done with a per-patch loss, $L_{\text{CNN}}$, which is a combination of weighted cross-entropy loss and dice loss. Weighted cross-entropy gives more importance to the rare class fire by increasing its contribution to the loss, while the dice loss measures overlap (i.e. intersection over union) and directly optimizes for better segmentation of small or imbalanced regions. The losses are as follows:
$L_{\text{CNN}}$ $=L_{\text{WCE}}+L_{\text{DICE}}$ (5)
$L_{\text{WCE}}$ $=-w_{\text{fire}}\sum_{i}y_{i}\log(\hat{y}_{i})-w_{\text{no-fire}}\sum_{i}(1-y_{i})\log(1-\hat{y}_{i})$ (6)
$L_{\text{DICE}}$ $=1-\frac{2\sum_{i}y_{i}\hat{y}_{i}}{\sum_{i}y_{i}+\sum_{i}\hat{y}_{i}}$ (7)
where $y_{i}$ is the ground truth label for a patch (1 for fire, 0 for no fire), $\hat{y}_{i}$ is the predicted probability for fire, and $w_{\text{fire}}$ and $w_{\text{no fire}}$ are class weights.
3.2 Transformer-based Architecture
In the three settings, our ViT architectures re-use most of the components of their CNN counterparts, as shown in Figure 8, 14, and 15, respectively. The main difference is the absence of multi-scale feature maps after the satellite image encoding in options [Uncaptioned image]satellite images only and [Uncaptioned image]satellite and environmental data, or after the mid-resolution encoding for option [Uncaptioned image]environmental predictors only.
For a given satellite image time series $x_{1:T}=\{x_{t}\}_{t=1}^{T}$, each image $x_{t}\in\mathbb{R}^{H\times W\times C}$ is encoded independently by a pre-trained ViT architecture, specifically DINOv2: ViT-S (Oquab et al., 2023): $f(x_{t})=\{z_{t}\}$, with $z_{t}\in\mathbb{R}^{H_{p}\times W_{p}\times D_{p}}$, which outputs one feature map per time-step. Similarly to the CNN architecture, the encoding of all satellite images time steps is done in parallel, and the feature maps are concatenated across the temporal axis: $z_{1:T}=\{z_{t}\}_{t=1}^{T}$. As for the CNN, the temporal encoding is also done via a ConvLSTM model: $g(z_{1:T})=s$, with $s\in\mathbb{R}^{H_{p}\times W_{p}\times D_{p}}$. Multi-scale feature maps are not necessary for ViT due to the native high-resolution of the final output feature map: $H_{p}$ and $W_{p}$. The output feature map, $s$, is interpolated to the label dimensions $H_{\text{fire}}=W_{\text{fire}}=\frac{H}{10}=\frac{W}{10}$ and finally passed to the model head, a patch classification layer, to output the class probabilities: $h_{\text{ViT}}(s)=\hat{y}\in[0,1]^{H_{\text{fire}}\times W_{\text{fire}}\times 2}$, with $h_{\text{ViT}}$ the function representing the interpolation, and classification layer.
In the multi-modal model ([Uncaptioned image]satellite and environmental data), the encoding of environmental inputs is identical to that of the CNN method. The final environmental encoded features $s_{\text{env}}\in\mathbb{R}^{D_{\text{env}}}$ is replicated spatially and concatenated with the final feature map $s$ before the patch classification layer to output the class probabilities: $h(s,s_{\text{env}})=\hat{y}\in[0,1]^{H_{fire}\times W_{fire}\times 2}$.
For option [Uncaptioned image]environmental predictors only, the same modifications from the satellite image time series are applied to the mid-resolution inputs; for low-resolution inputs, we use a ViT-S architecture similar to the one used for mid-resolution inputs, as it already represents the smallest available model for the DINOv2 architecture.
Contrary to the training for the CNN-based architectures, the loss used here is only the dice loss as defined in Equation 7, because experimentally it led to the best results.
Figure 8: ViT Architecture for Wildfire Prediction in setting Refer to captionsatellite and environmental data. Top: Satellite image time series encoding, Bottom: Environmental predictors encoding.
4 Results
This section details the key results for the benchmark models described in Section 3. CanadaFireSat covers the period 2016-2023: we train our models on the years 2016-2021, while keeping 2022 for validation (Val) and 2023 for both test sets: Test and Test Hard. Results are evaluated in terms of F1 score and PRAUC (Area Under the Precision-Recall Curve for the positive class fire only). Both metrics are robust to imbalanced datasets, contrarily to patch-level accuracy. The F1 score is defined as the harmonic mean between Precision (proportion of true positive pixels over pixels predicted as positive) and Recall (proportion of true positives over all actual positives):
$F1=2\times\frac{\textbf{Precision}\times\textbf{Recall}}{\textbf{Precision}+\textbf{Recall}}\,$, (8)
It provides information about how well the model minimizes both false negatives and false positives at a fixed threshold. We favor the F1 score over Intersection over Union (IoU) as the former is more commonly used in the wildfire forecasting literature. PRAUC summarizes the Precision-Recall trade-off across all thresholds for the class fire.
The benchmark models are tested against a baseline approach relying on the FWI in the following way: first, for a given time step $t$ we extract the 8-day mean FWI map at 0.250.25° from t8t-8 to t1t-1 included, and interpolate it at the target resolution of $100\>\text{m}\times 100\>\text{m}$, then, the per-patch prediction is obtained by binarizing the interpolated FWI map. The optimal threshold is tuned on the validation set, referring to the year 2022: $\text{FWI}_{\text{th}}=6$. The PRAUC is computed by scaling the FWI values with respect to the maximum value: $\text{FWI}_{\text{max}}=50$.
4.1 Performance Analysis
We evaluate the performance of the two different architectures (CNN and ViT) across three different settings described in Section 3: [Uncaptioned image]satellite images only, [Uncaptioned image]environmental predictors only, and [Uncaptioned image]satellite and environmental data. The results are reported in Table 4.
Encoder Modality Params (M) Val Test Test Hard Avg
PRAUC F1 PRAUC F1 PRAUC F1 PRAUC F1
ResNet-50 SITS Only 52.2 45.2 49.3 53.3 58.9 26.3 36.7 41.6 48.3
ENV Only 97.5 41.6 46.7 49.9 53.5 24.5 33.1 38.7 44.4
Multi-Modal 52.2 46.1 51.1 57.0 60.3 27.1 37.4 43.4 49.6
ViT-S SITS Only 36.5 45.2 50.6 51.2 51.9 25.7 33.8 40.7 45.2
ENV Only 54.8 34.8 45.7 49.2 59.9 21.2 35.1 35.1 46.9
Multi-Modal 37.7 43.9 50.0 56.3 59.2 25.1 36.6 41.8 48.6
Baseline (FWI) ENV Only - 20.0 32.7 43.1 50.3 21.1 32.7 28.1 38.6
Table 4: Performance comparison of different model settings. Bold indicates the best metric value for each dataset split and model type, and underline denotes the runner-up.
Across the three evaluation sets (last column of Table 4), both ResNet-50 and ViT-S trained on [Uncaptioned image]satellite and environmental data reach the highest performance, with $15\%$ PRAUC and $11\%$ F1 score for the CNN, compared to the FWI baseline. For both the CNN and ViT architectures, relying on multi-modal inputs shows, on average, an improvement over models trained on [Uncaptioned image]satellite images only or [Uncaptioned image]environmental predictors only. While individually [Uncaptioned image]satellite images only and [Uncaptioned image]environmental predictors only are already highly discriminative, the multi-modal setting [Uncaptioned image]satellite and environmental data remains the most accurate forecaster with an average gain of $1.8\%$ in PRAUC and $1.3\%$ in F1 score for CNN-based models and $1.1\%$ in PRAUC and $1.7\%$ in F1 score for the ViT. We further discuss the role of satellite image time series and environmental predictors in Section 5.1. On the Test set, the better performance for both F1 score and PRAUC of all models compared to the Val set can be explained by fire patterns likely being more easily distinguishable due to the extreme fire season, a behavior also observed in the FWI baseline. Nonetheless, our best performing CNN model relying on [Uncaptioned image]satellite and environmental data still outperforms the FWI baseline, on the Test set by $13.9\%$ (PRAUC) and by $10\%$ (F1 score). Further analysis of the model’s performance threshold analysis across the Val, Test, and Test Hard sets is shown in D. The drop in performance of all models on the Test Hard set demonstrates the impact of the sampling strategy and the necessity of such an evaluation set: Test Hard can be used to assess models’ lower bound performance and their ability to model the hidden phenomena behind ignition. When it comes to the comparison between ResNet-50 and ViT-S trained on [Uncaptioned image]satellite and environmental data, the former shows to perform best in terms of F1 score across all sets. However, differences remain small, and both architectures seem valid solutions for wildfire forecasting.
The performances of the different models in setting [Uncaptioned image]satellite and environmental data are studied in Figure 9 for increasing FWI values. We first focus on the False Positive Rate, defined as $\text{FPR}=\frac{\text{FP}}{\text{FP}+\text{TN}}$. As expected, we can observe in Figure 9(a) a positive correlation between the FPR and the FWI. Indeed, negative samples associated with a higher FWI show similar fire danger conditions to positive samples, and are thus much more difficult to discriminate, with ignition becoming the main triggering factor for samples with $\text{FWI}>20$. Then, we study in Figure 9(b) the variations of the weighted F1 score, defined as $\hat{\text{F1}}=\frac{\text{F1}-\text{F1}_{pos}}{1-\text{F1}_{pos}}$. This second index tells how good the model is compared to a naive predictor: $\text{F1}_{pos}$, assigning the class fire to all samples. We note a negative correlation between $\hat{\text{F1}}$ and the FWI: as the FWI increases, there is approximately no difference between our benchmark models and a naive predictor. Such behavior is not surprising as it is more likely for a sample associated to a high FWI to belong to the class fire, as confirmed by the increase in percentage of positive samples with higher FWI (from $13\%$ at $\text{FWI}\in[0,5]$ to $77\%$ at $\text{FWI}\in[20,30])$. We can conclude that the improved performance of our model with respect to the FWI baseline reported in Table 4 is due to a better prediction of wildfire occurrence at lower FWIs, as the task becomes trivial for $\text{FWI}>20$ due to data imbalance.
(a) Negative samples FPR for the Multi-Modal methods across FWI levels.
(b) Weighted F1 score for the Multi-Modal methods across FWI levels.
Figure 9: Refer to captionsatellite and environmental data models’ performance across different FWI value groups.
We also study the results of ResNet-50 and ViT-S trained on [Uncaptioned image]satellite and environmental data across the most common land cover classes in CanadaFireSat. Our models struggle the most on the classes wetland and cropland. Indeed, fire patterns in these two land cover types differ from those observed in the majority of wildfires, which tend to affect forest areas. In particular, peatland fires in Canada can occur under the ground in wet areas, or even under the snow layer. Such fires are difficult to observe through the predictors considered in the proposed CanadaFireSat, and would require ad-hoc modeling due to the specificities of such ecosystems. Low scores are also observed for cropland fires, which also present unique fire patterns, as the ignition is often human-induced and driven by a specific need for agricultural practices. As before, detecting these events seems hardly possible with our remote sensing-based system.
Figure 10: Refer to captionsatellite and environmental data models’ F1 score across the main land cover classes.
4.2 Deployment at Scale: Case Study
CanadaFireSat enables training deep models for high-resolution wildfire forecasting. As a result, our dataset makes it possible to deploy models capable of monitoring large regions at high-resolution. In this section, we demonstrate on a real use case how a model trained on our CanadaFireSat dataset could be deployed at a scale useful for wildfire management teams.
We chose as a case study a large wildfire that occurred in British Columbia on $2023/07/01$, illustrated in Figure 11. The first row displays the RGB composite of the region of interest of size $16\>\text{km}\times 22\>\text{km}$ acquired by Sentinel-2 right before the wildfire starts on $2023/06/06$. The fire scar polygons from NBAC are shown on the second row. The third row shows the binarized predictions of the CNN model trained in the multi-modal setting [Uncaptioned image]satellite and environmental data, on the positive samples overlapping with the considered ROI. We observe how well the model delineates the urban interface on the left side of the wildfire and the rough approximation of its boundaries on the right side of the fire. However, we can also see at the top of the Sentinel-2 image that the model overestimates the wildfire extent. This case study showcases the potential of CanadaFireSat to enable the deployment of models capable of monitoring large regions at the unprecedented resolution of $100$ m.
Original S-2 Tile
Target Fire Polygon
Model Predictions
Figure 11: Row 1 Sentinel-2 tile from $2023/06/06$ of size $16\>\text{km}\times 22\>\text{km}$ before a large wildfire in British Columbia. Row 2 Fire polygons for the large wildfire on $2023/07/01$ over the same tile. Row 3 Binary model predictions (in red) over the $2.64\>\text{km}\times 2.64\>\text{km}$ center-cropped positive samples (patches outlined in black).
5 Discussion
5.1 High-resolution Wildfire Forecasting via Multi-modal Learning
As previously demonstrated in (Pelletier et al., 2023; Chowdhury and Hassan, 2015; Yang et al., 2021), multi-spectral multi-temporal satellite data can be a valuable data source to forecast wildfires. Indeed, several spectral indices discriminative for wildfire forecasting can be extracted from Sentinel-2: normalized difference vegetation index (NDVI), normalized difference water index (NDWI), tasseled cap wetness, and channel histograms. The results reported in Table 4 demonstrate the potential of multi-spectral temporal satellite data for high-resolution wildfire forecasting (setting [Uncaptioned image]satellite images only). While hydrometeorological data (setting [Uncaptioned image]environmental predictors only) are commonly used in global and continental wildfire forecasting models, they can be complemented by satellite data to improve strongly both the spatial resolution and the accuracy of the prediction (setting [Uncaptioned image]satellite and environmental data), as reported in Table 4, where the multi-modal approach leads to the best performances.
5.2 Importance of Negative Sampling for Training and Evaluation
Numerous wildfire forecasting benchmarks require sampling the negative (non-fire) samples due to extreme imbalance and computational constraints. A common strategy is to focus on samples that burned once across the period studied (Bakke et al., 2023; Prapas et al., 2022, 2023). In CanadaFireSat, we opt for a different strategy: we sample negative examples for each Canadian province uniformly across their yearly fire driver patterns (FWI values). By sampling the training, validation, and test sets in this way, we aim to train and evaluate our models on a subset representative of the conditions encountered in all of Canada. Nonetheless, as the yearly fire patterns vary, the distribution of negatives with respect to the FWI changes over the years, in turn affecting the performance of models. This motivated the creation of a second test set to understand the impact of sampling on the models’ performance (Test Hard), where ignition, a complex phenomenon difficult to model (Chen et al., 2021; Calef et al., 2008), differentiates positive from negative patches. Indeed human-induced ignitions, generally caused by infrastructures, agricultural practices, or ”recreational” activities, are typically hard to estimate with CanadaFireSat, as Sentinel-2 is the only source providing information on human presence, but only for a limited spatial context of $2.64\>\text{km}\times 2.64\>\text{km}$. Fine-tuning our multi-modal models on data such as Test Hard or enhancing our set of predictors with proxies of ignition probability (e.g, proximity to human settlements, or lightning probability) are relevant directions for improving our models towards accounting for ignition probability.
5.3 Modeling Wildfires in the Boreal Ecosystem
As initially stated in Section 1, one of our main motivations is the rise of wildfires in the boreal ecosystem and the risks this represents for its local communities. To cover the areas of interest and to evaluate the broader impact of wildfires on global climate, we created our benchmark CanadaFireSat so that it covers the entirety of Canada, including all its agricultural lands, urban areas, and other ecosystems such as the temperate forest in British Columbia. To study the behavior of trained models on the boreal ecosystem, it is possible to constrain the analysis on the main land cover classes of the boreal ecosystem (needleleaf forest and wetlands). With our benchmark models trained on CanadaFireSat, we observe an important difference of performance between those two land cover classes: with Multi-Modal CNN and ViT performing respectively $11.9\%$ and $11.4\%$ better on needleleaf forest compared to wetlands in terms of weighted F1 score on both the Val and Test sets, showing that for the latter land cover, performance is still not optimal. Indeed, wetland wildfires are a unique phenomenon compared to forest wildfires, as they depend much more on soil-related predictors and can burn underground for a long period. As a consequence, they are sometimes undetectable for optical remote sensing satellites. In particular, peatland wildfires that emit large amounts of CO2 and mercury (Fraser et al., 2018; Kohlenberg et al., 2018) are commonly studied independently from forest fires (Pelletier et al., 2023; Bali et al., 2021). Extending CanadaFireSat so that it includes data acquired from radar remote sensing satellites (for instance, Sentinel-1 images) could help to better model the surface soil conditions for wetlands (Millard and Richardson, 2018) and bridge the gap in performance across the boreal ecosystem.
5.4 Operationalization of the Model
Deploying models trained with CanadaFireSat over the entirety of Canada would require densely sampling the country with Sentinel-2 image time series, resulting in a huge amount of data to be processed. Indeed, the proposed dataset is aimed at modeling wildfire patterns at a moderate scale, but at high-resolution, and can be coupled with coarser resolution approaches (Prapas et al., 2022; Bali et al., 2021) to identify areas of interest and then apply our model to map such areas more precisely. Such coupling would allow wildfire management experts to target specific areas at risk for fine-grained wildfire forecasting or focus on areas that require more surveillance due to their proximity to local communities or due to their ecological and environmental interest. By alleviating the need for significant computational resources, it would break the barrier to scale this approach to large continental areas. Learning models directly capable of multi-scale prediction is an interesting future research direction to deploy high-resolution wildfire forecasting at scale. Such a model would exploit hierarchical learning approaches developed in computer vision for semantic segmentation (Li et al., 2022; Atigh et al., 2022).
5.5 Limitations and Future Work
As mentioned in Section 5.2, the main limitation of methods trained on CanadaFireSat is the difficulty of modeling the ignition component in wildfires due to its inherent stochasticity. Weather data from ERA5 can provide information on the risk of lightning, nonetheless, explicitly adding lightning probability (Geng et al., 2019) as a predictor, as well as other proxies for human ignition like the proximity to human settlement could help the trained models to better characterize ignition.
Multi-task learning (Zhang and Yang, 2021) could also be leveraged to develop a model forecasting wildfires at multiple scales. One could leverage different forecasting heads at multiple resolutions: $10$ km, $1$ km, $100$ m. This could help alleviate memory size constraints when high-resolution forecasts are deemed unnecessary and help providing consistent predictions across scales.
Moreover, one could investigate the potential of geolocation embeddings such as SatCLIP (Klemmer et al., 2023) or GeoCLIP (Vivanco Cepeda et al., 2023) to represent high-resolution non-dynamic satellite information. These could be combined with non-spatial, but temporal dynamics from Sentinel-2 (Pelletier et al., 2023) as a way to factorize spatial and temporal components in satellite data and limit memory consumption. Extending CanadaFireSat with atmospherically corrected images (e.g. L2A) or with BRDF-corrected Harmonized Landsat and Sentinel-2 data could help improving performances.
Another line of future research deals with the improvement of the pretraining of our multi-modal deep learning approaches. In our work, we leverage image encoders pre-trained on natural images such as ImageNet or via DINOv2, which are very different from multi-spectral satellite images. With the drastic increase in availability of Earth observation data, several models are being proposed to learn in an unsupervised way generalizable representations from this data (Cong et al., 2022; Jakubik et al., 2023; Hong et al., 2024; Astruc et al., 2024; Sumbul et al., 2025). One could study the potential of those foundation models as pre-trained representations to be used in high-resolution wildfire forecasting; CanadaFireSat could be the perfect starting point for such an investigation.
Finally, the increased complexity of models raises concerns regarding their interpretability and the possibility of understanding the role of the input variables in the final predictions. Several approaches exist to provide interpretations of black box wildfire forecasting models via feature attributions (Sundararajan et al., 2017; Selvaraju et al., 2017) or ranking (Lundberg and Lee, 2017), or even to directly build interpretable wildfire forecasting model architectures (Koh et al., 2020; Chen et al., 2019) via dense prediction architecture (Sacha et al., 2023; Porta et al., 2025a). However, those methods need adaptation to accommodate multi-modal (Ekim and Schmitt, 2023; Wang et al., 2023) or multi-temporal data (Turbé et al., 2023; Gee et al., 2019; Ghosal and Abbasi-Asl, 2021). They are also often not directly applicable to Earth observation data (Porta et al., 2025b) due to their strong implicit bias for natural images (Chen et al., 2019). This gap remains unfulfilled, and future works, for and beyond the wildfires prediction problem, should explore interpretable methods specifically tailored to Earth observation problems.
6 Conclusion
In this paper, we introduced CanadaFireSat, a comprehensive benchmark dataset for high-resolution wildfire forecasting over Canada from 2016 to 2023. CanadaFireSat was constructed to support multiple settings for model training: [Uncaptioned image]satellite images only, [Uncaptioned image]environmental predictors only, and [Uncaptioned image]satellite and environmental data. We demonstrated experimentally the potential of multi-modal learning for high-resolution wildfire forecasting on CanadaFireSat across two architectures: ResNet and ViT. Moreover, our experiments showed the importance of negative sampling in the evaluation of wildfire forecasting models. CanadaFireSat aims to accelerate research towards high-resolution monitoring of at-risk regions of interest to support wildfire management teams who are tasked with monitoring and protecting vast areas, such as the boreal ecosystem covering much of Canada. Results from this work demonstrate the feasibility of constructing future datasets like CanadaFireSat for other fire-prone landscapes where high-resolution fire polygons are available, like the Pan-Arctic, Pan-boreal, and grassland and forest ecosystems of the Tropics, since all input variables are globally available and open-access, even though certain fire regimes might require other high-resolution sensors, as seen for peatland fires. We hope this dataset will foster research in this direction. To facilitate that, all codes, models, and CanadaFireSat are made publicly available on GitHub and HuggingFace.
Appendix A Model Architectures
In this section, we illustrate the architectures used in the settings [Uncaptioned image]satellite images only (Figure 12) and [Uncaptioned image]satellite and environmental data (Figure 13) for the CNN-based models. We then show those used in the Transformer-based models in Figures 14 and 15, respectively.
Figure 12: CNN Architecture for Wildfire Prediction used for setting Refer to captionsatellite images only.
Figure 13: CNN Architecture for Wildfire Prediction used for Setting Refer to captionenvironmental predictors only.
Figure 14: ViT Architecture for Wildfire Prediction used for Setting Refer to captionsatellite images only.
Figure 15: ViT Architecture for Wildfire Prediction used for Setting Refer to captionenvironmental predictors only.
Appendix B Experimental Setup:
B.1 CNN Architecture Parameters
As mentioned in Section 3.1, satellite image time series encoding is done via a ResNet-50 backbone pre-trained on ImageNet. During training, the inputs are of size $T=5$, $C=14$, and $H=W=240$ leading to a target resolution $H_{\text{fire}}=W_{\text{fire}}=24$. During testing, we compute the prediction on the whole sample of size $H=W=260$ with the same $T$ and $C$, leading to $H_{\text{fire}}=W_{\text{fire}}=26$. We extract the model’s last three feature maps of channel dimensions: 512, 1024, and 2048. Those feature maps pass through three independent ConvLSTM models, each one with kernel size $3\times 3$ and only one layer. The ConvLSTM models output feature maps of dimensions: 64, 128, and 256, which are then passed to a U-Net-like decoder and interpolated to the target size.
This model is extended to multi-modal inputs of dimensions $N_{\text{env}}=15$, including the day of the year, and $T_{\text{env}}=8$, as we leverage the whole time series for those inputs. This data is projected to a high-dimensional space of size $D_{\text{env}}=64$ and passed to an LSTM with one layer. The selected environmental predictors are the following: Total Precipitation Sum: 8-day Mean, Skin Temperature: 8-day Mean, Temperature (2m): 8-day Mean, Volumetric Soil Water Layer 1: 8-day Mean, Wind Speed (10m): 8-day Mean, Relative Humidity: 8-day Mean, Vapor Pressure Deficit: 8-day Mean, LST Day (1km): 8-day Mean, NDVI, EVI, FPAR, LAI, Drought Code: 8-day Mean, Fire Weather Index: 8-day Mean.
The model using only environmental predictors leverages inputs of dimension $H_{\text{mid}}=W_{\text{mid}}=32$ for mid-resolution data (MODIS), and $H_{\text{low}}=W_{\text{low}}=32$ for low-resolution data (ERA5, CEMS). MODIS data at $1$ km: LST Day is interpolated to $500$ m to align with the rest of the MODIS inputs. Similarly, for the CEMS data, the Fire Weather Index and Drought Code, originally at 0.250.25°, are interpolated to $11.1$ km to align with ERA5-Land. We leverage the same set of environmental predictors as in the multi-modal setting, split into the two resolution groups. The temporal dimension of those inputs is $T_{\text{env}}=8$, the number of mid-resolution predictors is $N_{\text{mid}}=6$ including the day of the year, and the number of low-resolution predictors is $N_{\text{low}}=10$ including the day of the year (one more dimension than $N_{\text{env}}$, as we include the day of the year twice). As mentioned in Section 3.1, for the mid-resolution group we leverage $N_{S}=5$ multi-scale feature maps of dimensions: $64$, $256$, $512$, $1024$, and $2048$, for the mid-resolution data. The last feature map of channel dimension $2048$ is one-dimensional and is passed to an LSTM network for the temporal encoding. For the other four, we use independent ConvLSTM models. Those temporal encoders output feature maps of dimensions $64$, $128$, $256$, $512$, and $1024$, which are passed to a U-Net-like decoder and interpolated to the target size. The low-resolution inputs are encoded via a smaller network: ResNet-18, which outputs feature maps of channel dimension $D_{\text{low}}=512$, encoded temporally with a LSTM with of one layer to a dimension $D^{\prime}_{\text{low}}=64$, matching the channel dimension of the last feature map of the U-Net decoder. Both ResNet encoders are pre-trained on ImageNet. As for the other settings, the training is done with a target resolution of size $H_{\text{fire}}=W_{\text{fire}}=24$, and at test time the target resolution is $H_{\text{fire}}=W_{\text{fire}}=26$.
B.2 ViT Architecture Parameters
In the case of the ViT architecture, the satellite image time series encoding is done via the DINOv2 ViT-S architecture. The input channel and temporal dimensions are the same as for the CNN architecture: $T=5$, and $C=14$. However, since the patch size of the ViT encoder is $14$, we used as input spatial dimensions a direct multiple: $H=W=252$ during training. As a consequence, during training $H_{\text{fire}}=W_{\text{fire}}=25$. At test time, input and target dimensions are the same as for the CNN use case. To reduce overfitting issues, we used the LORA method (Hu et al., 2022) to fine-tune the ViT model with rank $r=32$, $\alpha=32$, and dropout $d_{\text{LORA}}=0.1$. The channel dimension of the encoded feature map is $D_{p}=384$, which is maintained after temporal encoding via ConvLSTM with kernel size $3\times 3$.
The model extension for multi-modal data is done similarly to the CNN case with $D_{\text{env}}=384$. This is to match the channel dimension of the final feature map. The same set of environmental predictors is used for this setting as for the CNN architecture above.
For the environmental-only architecture, the input and target spatial dimensions and processing are identical to the CNN use case. The temporal dimension differs as we use $T_{\text{env}}=5$ for data augmentation. Both mid-resolution and low-resolution ViT-S encoders are randomly initialized and therefore do not use the LORA method for fine-tuning. In both encoders, for the position embedding, attention, and projection, we use a dropout rate $d_{\text{env}}=0.2$ and a stochastic depth rate of $d_{\text{depth}}=0.1$. For mid-resolution inputs, the patch size is $2$, and for low-resolution inputs, the patch size is $8$. The temporal encoding of the mid-resolution feature map is identical to the one used for the satellite image time series, and the temporal encoding of the low-resolution data is done through a one-layer LSTM with both input and output channel dimensions $D_{\text{low}}=384$.
B.3 CNN Training Parameters
The CNN models are trained using the combined weighted cross-entropy and dice loss. The positive class (fire) weight is $0.87$ and the negative class (no fire) is $0.13$, found experimentally. Training is run over $20$ epochs with a batch size of $24$ samples on a NVIDIA GeForce RTX 3080 Ti GPU. The scheduler for the learning rate follows a $2$-epoch warm-up from the starting learning rate of $1e^{-7}$ to the base learning rate of $5e^{-6}$. Then the learning rate follows a cosine annealing of one cycle to the minimum learning rate of $1e^{-7}$ over the rest of the epochs. The optimizer used is ADAMW with a weight decay of $0.01$. During training, the augmentation pipeline first randomly crops the satellite input images to the training resolution, then resizes the images with a scale $s\in[0.9,1]$. The images are randomly flipped horizontally and vertically, and Gaussian noise with variance $\sigma^{2}\in[0.01,0.1]$ is injected. Finally, we randomly sample the satellite image time series to extract $T=5$ images (or pad when necessary). At test time, we center-crop the images to the required resolution and select the last $T=5$ samples. For the multi-modal training, the non-spatial environmental data is not augmented, while for the environment-only architecture, we apply random horizontal and vertical flipping and Gaussian noise injection, similarly to the satellite image time series. The missing values in the environmental predictors, mainly caused by the NDVI and EVI as they are 16-day composites, are replaced during training with the value $0.0$.
B.4 ViT Training Parameters
Most of the ViT training parameters are the same as for the CNN models, except for the batch, which, despite also being $24$, is accumulated across two steps of $12$ for the ViT models. Moreover, during the training of the environmental-only use case, as we select $T_{\text{env}}=5$ time steps, it is also necessary to randomly sample across the $8$ available samples. Finally, at test time across all modalities, we use the native temporal length for each sample, $8$ for the environmental data, and a variable length for the satellite image time series. The processing of the missing values for the environmental predictors is the same as for the CNN-based architecture.
Appendix C Ablation Study of the Impact of Satellite Image Time Series
In Table 5, we analyze the performance of the multi-modal models in setting [Uncaptioned image]satellite and environmental data with respect to the usage of time series. We compare our full multi-modal model using satellite image time series against a version using only the most recent image available before the prediction. In practice, for the CNN-based model, this impacts the number of parameters in the U-Net decoder as $D_{i}>D^{\prime}_{i}$. Regardless of the architectures, the model performs best when presented with SITS rather than a single Sentinel-2 tile. As a consequence, we can hypothesize that dynamic factors directly linked to wildfire can be learned by the model from the temporal dimension of Sentinel-2.
Encoder SITS PRAUC F1
ResNet-50 42.4 48.3
ResNet-50 46.1 51.1
ViT-S 38.2 47.4
ViT-S 43.9 50.0
Table 5: Ablation study of SITS impact on the validation set performance.
Appendix D Test Hard Analysis
Figure 16 demonstrates the domain shift between the Val and the Test set, as the evolution of the F1 score with the probability threshold is centered around 0.5 for the Val set, presenting a normal behavior while being shifted towards a smaller threshold value for the Test set. As a consequence, the metrics in Table 4 might overestimate the model performance on the test set due to the extreme fire patterns during this year. For this purpose, we constructed the adversarial set named Test Hard for the year 2023 as described in Section 2.1.2. Figure 16 also shows the delta in performance between Test and Test Hard: the centering of the maximum value for Test Hard is closer to the 0.5 threshold, representing a better alignment with the model behavior on the Val set.
Figure 16: Analysis of the F1 score performance as a function of the probability threshold across all evaluation sets. The circle, square, and triangle represent the maximum value for each set.
Figure 17 presents the change in land cover distribution for the negative samples between the two sets, Test and Test Hard, with respect to the positive samples. The stratification sampling done in Test Hard better aligns the categorical distributions for the negative and positive populations.
(a) Land cover distribution for the Test set across positive and negative samples.
(b) Land cover distribution for the Test Hard set across positive and negative samples.
Figure 17: Comparison of the land cover distribution across the Test and Test Hard sets for the positive and negative samples.
Appendix E Deployment at Scale: Second Case Study
In Figure 18, we present another case study for our CNN-based model in setting [Uncaptioned image]satellite and environmental data. This example presents a large wildfire in Québec occurring on the $2023/07/05$, displayed over an RGB composite of a Sentinel-2 image of $14\;\text{km}\times 26\;\text{km}$. The predictions follow the same pattern as the actual wildfire, despite slightly overestimating its extent, as it can be seen on both sides of the Sentinel-2 tile, similarly to what we observed in Figure 11.
Original S-2 Tile
Target Fire Polygon
Model Predictions
Figure 18: Row 1 Sentinel-2 tile from $2023/06/28$ of size $14\>\text{km}\times 26\>\text{km}$ before a large wildfire in Québec. Row 2 Fire polygons for the large wildfire on $2023/07/05$ over the same tile. Row 3 Binary model predictions (in red) over the $2.64\>\text{km}\times 2.64\>\text{km}$ center-cropped positive samples (patches boundaries are outlined in black).